04. As noted in the Experiment 4 hand-out, the intensity I that results from superposition of two coherent light beams at the photodiode detector is given by I =i, +12 + 2112 cos ? , 2? where ? = k (sl-s2)-EAs is the phase difference in radians between the two beams due to having traveled different path lengths s, and s2. Defining r 2/, we can rewrite that as 12vT cos 8] When referring to a Michelson interferometer, this r factor reflects some imbalance between the two arms due to an imperfect beam splitter and/or optical losses somewhere. Now, when using a sodium lamp, the two wavelengths that make up the D line doublet are produced with somewhat different intensities but should have the same r factor due to the interferometer. Let the irradiance for the light of wavelength ?1 be 11, and that for wavelength ?2 be 1,-q11. Here, 11 and 12 are not the signals for the two arms, but simply the signals associated with the two different wavelengths. The relative strength of the signal:s traveling through the two arms is still r. a) (5 pts) Show that the net intensity measured by the photodiode can be written as 1 = 11 [(1 + r)(1 + q) + 2vr (cos ?? + q cos&J), where ?? and ?2 are associated with wavelengths ?1 and ?2. [Hint: think about how ?2 interferes with d2]
Answer #1

The intensity due to \lambda_1 is

i_1=I_1(1+r+2\sqrt{r}\cos \delta_1) ....(1)

The intensity due to \lambda_2 is

i_2=I_2(1+r+2\sqrt{r}\cos \delta_2) ....(2)

Note that the phase difference is different for the different wavelengths because the phase difference depends on wavelengths.

\delta_1=\frac{2\pi}{\lambda_1}\Delta s

and

\delta_2=\frac{2\pi}{\lambda_2}\Delta s

Now note that I_2 =q I_1

Therefore equation (2) becomes

i_2=qI_1(1+r+2\sqrt{r}\cos \delta_2)=qI_1(1+r)+2 q\sqrt{r}I_1\cos \delta_2

So the net intensity is

\begin{align} I&=i_1+ i_2=I_1(1+r+2\sqrt{r}\cos \delta_1)+qI_1(1+r)+2 qI_1\sqrt{r}\cos \delta_2 \nonumber \nonumber \\ \Rightarrow I &= I_1(1+r)(1+q)+2\sqrt{r}I_1(\cos\delta_1+q\cos \delta_2) \nonumber \end{align}

Hence the net intensity measured by the photodiode is

I = I_1\Big[(1+r)(1+q)+2\sqrt{r}(\cos\delta_1+q\cos \delta_2)\Big]